OpenStudy (rizags):

Please Help The sum of the first n terms in a sequence is always 1/n. Find the product of the first 2007 terms.

OpenStudy (rizags):

well, the sum of the first 2007 terms is 1/2007 right?

OpenStudy (mathmath333):

if i write \(\dfrac{1}{4}\) ,\(2\) times then \(\dfrac{1}{4}+\dfrac{1}{4}=\dfrac{1}{2}\) or if i write \(\dfrac{1}{9}\) ,\(3\) times then \(\dfrac{1}{9}+\dfrac{1}{9}+\dfrac{1}{9}=\dfrac{1}{3}\) or if i write \(\dfrac{1}{16}\) ,\(4\) times then \(\dfrac{1}{16}+\dfrac{1}{16}+\dfrac{1}{16}+\dfrac{1}{16}=\dfrac{1}{4}\)

OpenStudy (rizags):

therefore, you are saying that the sequence might be \[\frac{ 1 }{ 2007^{2} }\times2007\]

OpenStudy (mathmath333):

so if i will write \(\Large (\dfrac{1}{n^2}) \) , \(\Large n\) times then then will be equal to \(\Large \dfrac{1}{n}\)

OpenStudy (mathmath333):

for the sequence u are given the term \(\Large\dfrac{1}{n^2}\)

OpenStudy (mathmath333):

let me call some this seems complicated @ganeshie8 ,@freckles

OpenStudy (mathmath333):

@freckles

OpenStudy (asnaseer):

The first term, \(a_1\) must be 1 The second term is given by: \(a_1+a_2=\frac{1}{2}\), so \(a_2=\frac{1}{2}-a_1\) The third term is given by: \(a_1+a_2+a_3=\frac{1}{3}\), therefore \(a_3=\frac{1}{3}-(a_1+a_2)=\frac{1}{3}-\frac{1}{2}\) So, in general: \(a_i=\frac{1}{i}-\frac{1}{i-1}=-\frac{1}{i(i-1)}\)

OpenStudy (asnaseer):

The product is therefore given by:\[P_n=1\times\left(-\frac{1}{2.1}\right)\times\left(-\frac{1}{3.2}\right)\times\left(-\frac{1}{4.3}\right)...\times\left(-\frac{1}{n(n-1)}\right)\]

OpenStudy (asnaseer):

can you spot a pattern in this?

OpenStudy (asnaseer):

HINT: Look for factorials in the denominator

OpenStudy (rizags):

ummm..... i see a pattern but im not sure how to put it into words

OpenStudy (asnaseer):

Ignoring the signs for now, can you see this: